OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 4) = 2x. 2 -2 -8 -4

OpenStudy (jamanorthi):

put on the numbers on x where it should and solve it

OpenStudy (anonymous):

can u show me on an equation or drawing please

OpenStudy (jamanorthi):

if x=2 then the equation will be 3(2-4)=2(2)

OpenStudy (anonymous):

okay then we distribute the three right ?

OpenStudy (jamanorthi):

yeah

OpenStudy (jamanorthi):

do the same thing with all the different number for x that is given there

OpenStudy (anonymous):

okay so (6-8) = ( 4)?

OpenStudy (jamanorthi):

(6-8)=4 3=4

OpenStudy (anonymous):

cool so the answer is -4 or no ( out of the choices)

OpenStudy (jamanorthi):

i am sorry i cannot give you an answer

OpenStudy (anonymous):

oh darn :(

OpenStudy (jamanorthi):

sorry :( bye